OpenStudy (anonymous):

Find the product of the complex numbers and enter it below. Remember that i = . (4 - i )(1 + 4i )

OpenStudy (neonumbrella5115):

Have you tried FOILing?

OpenStudy (anonymous):

no i haven't @neonumbrella5115 i dont know how to do this at all

OpenStudy (neonumbrella5115):

Okay so to FOIL you would multiply each term in the first parenthesis with both terms in the second parentheses. Then you would multiply the second term in the first parentheses by each term in the second parentheses. The add all those values. Example will follow.

OpenStudy (anonymous):

\[\left( 4-i \right)\left( 1+4i \right) =(4+15i-4i^2)\]

OpenStudy (triciaal):

use the distributive property twice also commonly referred to as FOIL because of the steps in the operation for complex numbers remember that i^2 = -1

OpenStudy (anonymous):

now you know that i=squrt(-1) and i^2=-1

OpenStudy (neonumbrella5115):

4*1=4 4*4i=16i -1i*1=-i -i*4i=-4i^2 Add those numbers together now. Then because i^2 is 1, plug that in as i^2

OpenStudy (triciaal):